OpenStudy (anonymous):

Simplify completely sqrt12+sqrt50-sqrt48

OpenStudy (amistre64):

might wanna think about seeing if there are any perfect squares hiding in the factors

OpenStudy (anonymous):

50 ;O ?

OpenStudy (amistre64):

... not quite if you list the perfect squares (1^2, 2^2, 3^2, 4^2, ...) its easier to see if they can factor into those things

OpenStudy (amistre64):

1^2 = 1 ; that ones a pointless one 2^2 = 4 3^2 = 9 4^2 = 16 5^2 = 25 those should give us enough to work with

OpenStudy (amistre64):

now take sqrt(12) can 12 be factored into a perfect square times something else?

OpenStudy (anonymous):

uhhhh 12 ? to 144

OpenStudy (amistre64):

12^2 does equal 144, but my thought prolly wasnt clear enough can we take any of the perfect squares from the list, and make 12?

OpenStudy (amistre64):

4*x = 12 16*x=12 25*x=12 like this

OpenStudy (anonymous):

4x3 ?

OpenStudy (amistre64):

good; lets use that to simplify; sqrt(12) sqrt(12) = sqrt(4x3) sqrt(4x3) = sqrt(4) x sqrt(3) sqrt(4) x sqrt(3) = 2sqrt(3) and thats as far as we can go with that one notice that finding a perfect square factor allowed us to move it to the ouside in the end

OpenStudy (amistre64):

can you try out the same concept with: sqrt(50)

OpenStudy (anonymous):

25 right ?

OpenStudy (amistre64):

yes, 25x2 sqrt(25x2) = sqrt(25) x sqrt(2) sqrt(25) x sqrt(2) = 5sqrt(2) very good this gets us to 2sqrt(3) + 5sqrt(2) - sqrt(48) we just have the 48 part to deal with now

OpenStudy (anonymous):

uhh 24 ?

OpenStudy (amistre64):

24 aint a perfect square, so its not going to help us out :)

OpenStudy (anonymous):

oh so its 5sqrt2-2sqrt3

OpenStudy (amistre64):

remarkably yes

OpenStudy (anonymous):

awwe ^,^ ! YES! thank you !

OpenStudy (amistre64):

youre welcome :) and good luck